OpenStudy (anonymous):

Where are the asymptotes of f(x) = tan 2x from x = 0 to x = π?

OpenStudy (anonymous):

x = 0, x = π, x = 2π x = 0, x = pi over 4 x = pi over 4, x = 3 pi over 4 x = pi over 2, x = 3 pi over 2 Those are the options... i think its b or c

OpenStudy (sidsiddhartha):

to calculate asymptotes of a given function u have to equate denominator with zero

OpenStudy (anonymous):

yah thats what i did but i got pi/4

OpenStudy (anonymous):

would it also be 0?

OpenStudy (aum):

tan(x) is undefined when x = pi/2 and x = 3pi/2 for 0 < x < 2pi Therefore, tan(2x) is undefined when x = pi/4 and x = 3pi/4 for 0 < x < pi

OpenStudy (sidsiddhartha):

vertical asymptotes are the points at which the function is undefined

OpenStudy (anonymous):

so pi/4 and 0 or pi/4 and 3pi/4

OpenStudy (aum):

tan(x) = sin(x) / cos(x) cos(x) = 0 when x = pi/2, 3pi/2 in the interval [0, 2pi] Therefore, tan(x) has a vertical asymptote at x = pi/2, 3pi/2 in the interval [0, 2pi] tan(2x) will have a vertical asymptote when 2x = pi/2, 3pi/2 or when x = pi/4, 3pi/4 in the interval [0, pi]

OpenStudy (anonymous):

so d?

OpenStudy (aum):

x = pi/4, 3pi/4

OpenStudy (anonymous):

oh ok thanks

OpenStudy (aum):

yw. (It is C)

OpenStudy (anonymous):

could you help with more? ill best answer and post them